OpenStudy (anonymous):

A rectangle box with a square base contains 8 cubic feet. If the height is 6 inches, how many feet are there in each side of the base?

OpenStudy (anonymous):

Let x = the side of the base , " with a square base" that's mean that x^2 = area of the base .. you 're given the height as 6inch convert it into ft and it will be 5 ft . THERFORE , .5x^2 = volume of the box=8 5x^2=8 Solve for x .

OpenStudy (anonymous):

Got it ?

sam (.sam.):

If it's a square base how can it be 8 'cubic' feet?

OpenStudy (anonymous):

Answer : 4, right?